# DuoSpaceNet: Leveraging Both Bird’s-Eye-View and Perspective View Representations for 3D Object Detection

### v1

@@ 4.2 Map Segmentation Results
In Tab. 2, we benchmark the map segmentation performance on nuScenes val set. All methods use ResNet-101-DCN backbone except for M2BEV, who has a more advanced backbone. We only carry out single-frame experiments due to limited GPU availability.
Compared with previous single-frame methods, our model achieves the highest IoU for both $drivable\ area$ and $lane\ boundary$, regardless of whether the segmentation branch is trained jointly with object detection or not. When it comes to the multi-frame setting, our single-frame model outperforms most previous models such as BEVFormer and UniAD [8] by a large margin. Compared with the current leading temporal method PETRv2, our single-frame model still excels in lane segmentation (46.5 IoU v.s. 44.8 IoU) and achieves comparable performance in drivable area (81.2 IoU v.s. 83.8 IoU).
@@ 4.3 Ablation Studies
To make BEV features more distinctive from PV features, we propose adding feature divergence enhancement during BEV feature generation. As shown in Tab. 5, while adding it to the BEV-only baseline can improve mAP by 0.7%, it won’t yield any help on NDS. Adding feature divergence enhancement in conjunction with our duo space design, however, will significantly improve the mAP by 1.3% and NDS by 0.6%, benefiting from the contrastiveness added between BEV and PV features.
Table 6: Ablation of the proposed Duo Space Queries.
@@
Duo Space Queries.
Although using feature maps from both spaces inherently has advantages over from single space, optimal performance cannot be achieved without our delicately designed duo space object queries. To validate this, three models differing only in their decoders were obtained. The first model, “unshared pose and unshared content”, divides classical object queries into two sets, each attending separately to either BEV or PV features in cross-attention layers. The second model, “shared pose and shared content”, makes each classical object query sequentially pass through self-attention, PV and BEV cross-attention layers, thus sharing pose and content embedding across both spaces. The third model,“shared pose and unshared content”, is equipped with our proposed duo space object queries. As Tab. 6 reveals, the first setting yields the worst result. In such a setting, each query only sees a single space of features, thereby losing the potential advantages offered by the presence of duo space features. The performance is marginally improved when pose and content embeddings are both shared, while the best results are achieved coupled with our Duo Space Decoder design. This implies that directly sharing content features across spaces might create latent space confusion due to different feature distributions. Decoupling content embeddings, however, proves to be more optimal by preserving unique feature representations from both spaces.
Figure 5: Visualization of 3D detection results in top-down view. Ground truth bounding boxes are in green and predictions are in blue
Figure 6: Visualization of 3D detection results in perspective camera view. Different colors represent different categories.
Table 7: Ablation of different temporal strategies.
@@
We demonstrate the necessity of a unified temporal solution for both spaces in contrast to some trivial solutions. We keep using temporal queries in PV space across all experiments. In each experiment, we use a different temporal strategy in BEV space. Specifically, “Recurrent Stacking” refers to infusing temporal information by stacking up temporal BEV features. “Learnable Attention” refers to infusing temporal information by temporal self-attention proposed in BEVFormer. “Temporal Queries” refers to our method where both spaces infuse temporal information into their temporal duo space queries. As clearly shown in Tab. 7, temporal strategy matters a lot. The proposed Duo Space Temporal Modeling achieves far superior performance compared with simply using an off-the-shelf popular BEV temporal method in our duo space paradigm.
@@ 4.4 Qualitative Analysis
In addition to the quantitative analysis, we report qualitative results of our DuoSpaceNet in comparison with PV-only and BEV-only baselines. In this example, as top-down-view visualization (Fig. 5) reveals, the bounding boxes (highlighted by the orange ellipse) predicted by DuoSpaceNet align more closely with the ground truth in 3D space compared to the PV-only baseline. This alignment is attributed to the more explicit 3D spatial information in BEV features. Regarding the object pose estimation (the red circle in Fig. 5, depicting a motorcycle in Fig. 6), due to the significant appearance information in PV features, both DuoSpaceNet and the PV-only baseline accurately predict its heading angle. In contrast, the BEV-only baseline provides an erroneous prediction differing by 90 degrees from the ground truth (Fig. 5 middle column). Combined, we are confident that our duo space paradigm achieves “the best of both world”.
5 Conclusion
In this paper, we introduce DuoSpaceNet, a camera-based 3D perception framework that aggregates information from both BEV and PV spaces. With our proposed duo space decoder design, our model is able to benefit from dense BEV representations and lossless PV features at the same time and achieve better performance on 3D object detection and map segmentation tasks. Ablation studies demonstrate the effectiveness of each individual design we propose, and extensive experiments on nuScene dataset verify that our proposed DuoSpaceNet achieves competitive performance against state-of-the-art methods.
@@ 5 Conclusion
We then generate $\mathbf{z}_{BEV}^{(t)}$, $\mathbf{z}_{PV}^{(t)}$, $\hat{\mathbf{p}}_{BEV}^{(t)}$ and $\hat{\mathbf{p}}_{PV}^{(t)}$, $t\in\{T-l+1,T-l+2,...,T\}$ according to Eq. 2368 as temporal inputs. Finally, after cross-attention layers (Eq. 7 and 9), we aggregate temporal outputs via 3-layer MLP before the FFN of each decoder layer. An illustration of temporal cross-attention in PV space is shown in Figure 7. The temporal cross-attention in BEV space is identical expect for the use of BEV space queries and BEV features as input.
@@
8 Future works.
We identify some interesting topics for future exploration using our duo space framework. First, our method enables some creative ways of data augmentation. For example, during training time, we can randomly mask out the entire BEV or PV feature maps, making the model more robust to feature variations across different spaces.
Second, our framework is capable of long-range 3D detection. Since rays captured by camera lens originated from a distant position cannot be fully reconstructed within a fixed-size BEV feature map, a cut-off perception range and quantization are applied during BEV representation generation to achieve computational efficiency in practice. Consequently, distant objects are inevitably lost during this view transformation process. Our framework, however, is capable of detecting objects that are out of the range of our BEV feature map by looking into their PV features. We argue that our method is more aligned with the real-world needs of self-driving. We are excited to benchmark its long-range performance if a proper dataset is available in the future.
Third, we are aware that due to the nature of 3D convolution, the feature divergence enhancement module we proposed introduces a significant number of computations. While the use of the enhancement module is complementary (i.e., without which our duo space design still works), one way to mitigate its computational burden is to use LSS-like [30] 2D to 3D feature lifting methods in conjunction with sparse convolution [4]. This approach can also enhance the contrastiveness between two spaces with less computations.

### v3

@@ 4.2 Map Segmentation Results
In Tab. 3, we benchmark the map segmentation performance on nuScenes val set. All methods use ResNet-101-DCN backbone except for M2BEV, who has a more advanced backbone. Compared with previous methods, our model achieves the highest IoU for both $drivable\ area$ and $lane\ boundary$, regardless of whether the segmentation branch is trained jointly with object detection or not. While sharing modules between different tasks saves computational resources and reduces inference time, the jointly trained model does show a slight performance drop compared to the individually trained model on map segmentation. This phenomenon, known as the negative transfer effect, is consistent with findings from previous studies [18].
@@ 4.3 Ablation Studies
To make BEV features more distinctive from PV features, we propose adding feature divergence enhancement (FDE) during BEV feature generation. As shown in Tab. 6, while adding it to the BEV-only baseline can improve mAP by 0.7%, it won’t yield any help on NDS. Adding FDE in conjunction with our duo space design, however, will significantly improve the mAP by 1.3% and NDS by 0.6%, benefiting from the contrastiveness added between BEV and PV features.
Figure 5: Qualitative comparison of top-down 3D detection results among our method, PV-only, and BEV-only models. Ground truth bounding boxes are in green and predictions are in blue. Our prediction aligns most accurately with the ground truth. Please refer to the supplementary materials for more visualization.
@@
Table 7: Ablation of the proposed Duo Space Queries.
@@
Table 8: Ablation of different temporal strategies.
Duo Space Queries.
Although using feature maps from both spaces inherently has advantages over using those from a single space, optimal performance cannot be achieved without our delicately designed duo space object queries. To validate this, three models differing only in their decoders were obtained. The first model, “unshared pose and unshared content”, divides classical object queries into two sets, each attending separately to either BEV or PV features in cross-attention layers. The second model, “shared pose and shared content”, makes each classical object query sequentially pass through self-attention, PV and BEV cross-attention layers, thus sharing pose and content embedding across both spaces. The third model,“shared pose and unshared content”, is equipped with our proposed duo space object queries. As Tab. 7 reveals, the first setting yields the worst result. In such a setting, each query only sees a single space of features, thereby losing the potential advantages offered by the presence of duo space features. The performance is marginally improved when pose and content embeddings are both shared, while the best results are achieved coupled with our Duo Space Decoder design. This implies that directly sharing content features across spaces might create latent space confusion due to different feature distributions. Decoupling content embeddings, however, proves to be more optimal by preserving unique feature representations from both spaces.
@@
We demonstrate the necessity of a unified temporal solution for both spaces in contrast to some trivial solutions. We keep using temporal queries in PV space across all experiments. In each experiment, we use a different temporal strategy in BEV space. Specifically, “Recurrent Stacking” refers to infusing temporal information by stacking up temporal BEV features. “Learnable Attention” refers to infusing temporal information by temporal self-attention proposed in BEVFormer. “Temporal Queries” refers to our method where both spaces infuse temporal information into their temporal duo space queries. As clearly shown in Tab. 8, temporal strategy matters a lot. The proposed Duo Space Temporal Modeling achieves far superior performance compared with simply using other popular designs.
@@ 4.4 Qualitative Analysis
In addition to the quantitative analysis, we report qualitative results of our DuoSpaceNet in comparison with PV-only and BEV-only baselines. In this example, as top-down-view visualization (Fig. 5) reveals, the bounding boxes (highlighted by the orange ellipse) predicted by DuoSpaceNet align more closely with the ground truth in 3D space compared to the PV-only baseline. This alignment is attributed to the more explicit 3D spatial information in BEV features. Regarding the object pose estimation (the red circle in Fig. 5, depicting a motorcycle in the real world), due to the significant appearance information in PV features, both DuoSpaceNet and the PV-only baseline accurately predict its heading angle. In contrast, the BEV-only baseline provides an erroneous prediction differing by 90 degrees from the ground truth. Combined, we are confident that our duo space paradigm achieves “the best of both world”.
5 Conclusion and Limitations
In this paper, we introduced DuoSpaceNet, a camera-based 3D perception framework that effectively integrates features from both BEV and PV spaces. The proposed duo-space decoder leverages dense BEV representations alongside detailed PV features, significantly enhancing performance on 3D object detection and map segmentation tasks. Extensive experiments on the nuScenes dataset demonstrate that DuoSpaceNet achieves competitive performance compared to state-of-the-art approaches, and ablation studies confirm the effectiveness of each proposed design. Our method can also benefit end-to-end self-driving [12, 43]. Beyond autonomous driving, our method holds potential for applications in human-centric tasks, such as human pose estimation [51, 52] and emotion understanding [44, 53].
Although our framework theoretically supports long-range 3D detection better than BEV-only methods, full evaluation is limited by insufficient far-range annotations in the nuScenes dataset. Future work will include comprehensive evaluations on larger datasets like Argoverse2 [50]. Additionally, current map segmentation relies exclusively on BEV features; integrating PV features could further enhance accuracy, particularly for detailed map structures.
@@ 5 Conclusion and Limitations
We then generate $\mathbf{z}_{BEV}^{(t)}$, $\mathbf{z}_{PV}^{(t)}$, $\hat{\mathbf{p}}_{BEV}^{(t)}$ and $\hat{\mathbf{p}}_{PV}^{(t)}$, $t\in\{T-l+1,T-l+2,...,T\}$ according to Eq. 2368 as temporal inputs. Finally, after cross-attention layers (Eq. 7 and 9), we aggregate temporal outputs via 3-layer MLP before the FFN of each decoder layer. An illustration of temporal cross-attention in PV space is shown in Fig. 6. The temporal cross-attention in BEV space is identical expect for the use of BEV space queries and BEV features as input.
@@
Figure 7: Visualization of 3D detection results in perspective camera view. Different colors represent different categories. Our method achieves the best prediction result for the motorcycle instance w.r.t. its 3D position as well as its orientation, showcasing the effectiveness of incorporating both PV and BEV information in detection queries.
8 Additional Visualizations
In Fig. 7, the 3D detection results are displayed in the perspective camera view for the same example as shown in Fig. 5, comparing three different detection methods along with the ground truth.